OpenStudy (anonymous):

If b^a = 2a + b , then find 2^3 + 3^2 +1?

OpenStudy (anonymous):

\[2\text{^}3 = 2(3) + 2\]\[3\text{^}2 = 2(2) + 3\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

No problem.

OpenStudy (anonymous):

2^3=2(3)+2 = 6 + 2 = 8 3^2=2(2)+3 = 4 + 3 = 7 1 = 1^1 = 2(1) + 1 = 2 + 1 = 3 Hence, 2^3 + 3^2 +1= 8 + 7 + 3 = 18